OpenStudy (anonymous):

PLZ HELP WILL MEDAL + FAN!!!! Which fraction expresses ��������� 16/28 in lowest terms? A. ��������� - 8/14 B. ��������� - 4/7 C. ��������� -4/14 D. - 2/7

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Divide the numerator and the denominator by 2

OpenStudy (anonymous):

No direct answers @Brostep0s

OpenStudy (anonymous):

B

OpenStudy (anonymous):

My apologies @Night-Watcher

OpenStudy (anonymous):

ok i will @Night-Watcher

OpenStudy (anonymous):

and @Brostep0s it is not B. but thank you all XD

OpenStudy (anonymous):

Just a note for the future @Brostep0s :)

OpenStudy (anonymous):

How isn't it?

OpenStudy (anonymous):

if u do wat @Night-Watcher said and divide the numrator and denomator by 2

OpenStudy (anonymous):

Once you divide them both by 2, do it again. And repeat this until you can no longer divide these numbers.

OpenStudy (anonymous):

ok wat if u have to find a equvilent fraction for this fraction : \[\frac{ 3 }{ 7}\]

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

Answer Choices?

OpenStudy (anonymous):

If not, you just have to multiple the numerator and the denominator by the same number (It could be any number). Once you do the result would be equivalent.

OpenStudy (anonymous):

o yey srry LOL hear: A \[\frac{ 15 }{ 35 }\] B \[\frac{ 6 }{ 10 }\] C\[\frac{ 9 }{ 14 }\] D. \[\frac{ 21 }{ 42 }\] @Night-Watcher

OpenStudy (anonymous):

ok i got the answer thank you

OpenStudy (anonymous):

No problem.